ShadowKid3:

Find the percent change: $0.25 (old) $0.36 (new) Please explain the answer to me

Tbone:

Hope that helped

Dolphan:

Solution: Calculate percentage change from V1 = 0.25 to V2 = 0.36 (V2−V1)|V1|×100 =(0.36−0.25)|0.25|×100 =0.110.25×100 =0.44×100 =44%change =44%increase

Dolphan:

I think that is it. I don't know for sure.